OpenStudy (testing123guest):

Which number equals (3)^-3?

OpenStudy (calculusxy):

so the rule is: \(\large x^{-1} = \frac{1}{x^1}\)

OpenStudy (calculusxy):

this is the same thing really: \(\large 3^{-3} = \frac{1}{3^3} = \frac{1}{9}\)

OpenStudy (calculusxy):

sorry \(\frac{1}{27}\)

OpenStudy (testing123guest):

is that the answer

OpenStudy (calculusxy):

i think so .

OpenStudy (testing123guest):

cool i will let u know later if it is can u help me with a few more plz

OpenStudy (calculusxy):

i need to go and eat lunch.. sorry :(

OpenStudy (testing123guest):

cool peace dude